answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

If I marry someone with a valid work permit what are the chances that he'll be sent back to his country?

Now, I may be incorrect, but I am fairly confident that if you marry someone who is not a legal citizen of the country, and YOU are a legal citizen of the country, then he will be issued a green card, which grants him permanent resident status.

Probability to get exactly 2 heads when flipping a coin 3 times knowing that you get at least one head?

Okay, lets write out the possible outcomes when flipping a coin 3 times:

HHH, HHT, HTH, THH, TTH,THT,HTT,TTT

That constitures 8 scenarios in which the coin can fall over a 3 flip trial.

Now, it is known that you got "at least one head" so therefore we can rule out the no head scenario (TTT) which leaves us with 7.

Of those 7 times, how many times does it fall heads exactly twice?

Well, we have HHT,HTH,THH.

From this you can say that it there are 3 possible outcomes in which you get exactly two heads given that you get at least one head. 3/7.

If the probability of getting identical is 0.004 while the probability of fraternal twins is 0.023 then what is the probability that a person chosen at random will not be a twin?

1- P(identical) - P(fraternal) =1-0.004-0.023 =0.973 The probability of being a identical or fraternal twin plus the probability of not being a twin has to add to 1. so 1- probability of being twins=probability of not being a twin ;-)

What does perpendicular lines look like?

They are straight lines that intersect each other at right angles or 90 degrees.

What is tossing off?

basiclly you do the same what a boy does wanking, i only ever did it once, but they sed it felt auesome, tbh it seemed he like it better when i did it fast, but i sed it was my first time so i needed help, you basiclly have to put youre hand around there penis but not hard or tight just normally and then just move it with like youre wriest. im only very young, only a teenager, but we have to try these thnigs, + im sure practice makes it better :) xx

How do you find the probability of an event?

There are two main methods: the empirical method and the theoretical method.

Suppose you want the probability of throwing a 6 on one roll of a die.

Empirical method: Throw the die lots and lots of times and count the total number of throws as well as the number of throws that result in a 6. The probability of throwing a six is the number of times you get a 6 divided by the total number of throws.

Theoretical method: A die is a cube with six equal faces. Assuming the die is fair, each face has an equal probability of turning up. There are no other outcomes possible (the probability of the die ending up on an edge or a vertex are assumed to be zero). So, there are six outcome, and each equally probable, so that each has a probability of 1/6.

How does probability work?

Probability is the likelihood something will happen as opposed to another factor. If you were to toss a coin, the probability of it landing on heads is 1/2. If you toss it again, the probability of getting tails is 1/2.

Sometimes one result can affect the next one, other times it cannot,

The result affects the next if: You have 3 red marbles and 3 blue. You pull out a blue one and do not replace it. It affects is as it could not be chosen again so the probability of another blue goes down 1, as well as the mail number.

It would not if: You do the same as above, but DO replace the removed marble. This is because it can be chosen again, and the Probability and the overall amount stays the same.

A drawer cotains 50 bolts and 150 nuts. Half the of the bolts are rusted if one item is chosen at random what is the probability that it is rusted or a bolt?

P(A) chance of being a bolt = 50/200=1/4

P(B) i.e chance of a rusted item ( i.e bolt or nuts) =100/200=1/2

P(AB) = P ( OF BEING RUSTED AND BOLT)= 25/200= 1/8

P(A or B) = the probability that it is rusted or a bolt. = P(A) + P(B) -P(AB)=0.25+0.5-0.125=0.625

51/200

If two people are selected at random what is the probability that they were both born in may?

assuming that the probability of being born in may is equal to all other months then one person born in may is 1/12, so 2 people born in may is 1/12*1/12=1/144

What is the probability to obtain in sum the numbers 7 and 13 by throwing 3 dice?

Probability(Total = 7) = 15/216 = 0.069

Probability(Total = 13) = 21/216 = 0.097

A test has 2 multiple choice questions each with 5 choices what is the probability of guessing the correct answers to both questions?

You have a 4 percent chance of guessing both answers correctly assuming there is only one correct answer to each question and that you may only answer once per question.

What is population mean in statistics?

The population mean is the mean value of the entire population.

Contrast this with sample mean, which is the mean value of a sample of the population.

What is the probability the asteroid 99942 Apophis will strike earth?

The answer to this can be found in the links I will place below.

If you toss 4 coins what is the probability of getting two heads and two tails?

When tossing a coin, there are 2 distinct possibilities: heads or tails. When tossing 4 coins, there are 2 x 2 x 2 x 2 = 16 possible outcomes. To determine the number of outcomes that will result in 2 heads and 2 tails the formula would be n!/(h!)(n-h)! where n = number of coinsand h = number of heads. In this case the number of different ways you can get 2 heads and 2 tails would be

4!/(2! x 2!) = 6 Let's verify the number of different ways you can get two heads and two tails: H H T T

H T T H

H T H T T T H H

T H H T

T H T H These are the six different ways to get two heads and two tails with four coins. To find the probability, you divide 6 by the total number of possible outcomes (16) and you would get 6/16 = 3/8. The probability of tossing 4 coins and getting two heads and two tails is 3/8 or 0.375